The Global Oscillation Network Group (GONG) Project is a community-based activity to operate a six-site helioseismic observing network, to do the basic data reduction and provide the data and software tools to the community, and to coordinate analysis of the rich data set that is resulting. GONG data are available to any qualified investigator whose proposal has been accepted; however, active membership in a GONG Scientific Team encourages early access to the data and the collaborative scientific analysis that the Teams are undertaking. Information on the status of the Project, the scientific investigations, and access to the data are available on our WWW server whose URL is www.gong.noao.edu.
We are happy to report that the GONG network has provided excellent coverage of the Sun during the third quarter of 1997. The GONG instrument continues to operate at a high rate of reliability at all six sites. As reported for prior observing periods, much of the down-time occurred either at night or during periods of poor weather. Lost images due to instrumental problems remain at a minimum, with many of the image gaps filled by the simultaneously observed data from adjacent sites.
Down-time was associated with unusually severe winter weather at the Cerro Tololo site in Chile. Although no damage was incurred by the GONG instrument during the recent earthquake (with the exception of a blown fuse), communications to the site were lost for several days. The GONG instrument remained down for over 90 hours, but the event was overshadowed by the severity of damage sustained by the observatory and its infrastructure. Generous thanks to the staff at CTIO. In the midst of the extraordinary number of other problems facing them at the time, they were able to help track down the problem and bring the GONG instrument back on line.
Other weather-related down-time occurred at the Hawaii site when the Mauna Loa instrument suffered damage during a lightning storm. Communications were lost, other instruments around the observatory were affected, and some damage was incurred by the GONG electronics. The instrument was down nearly six days before the problems were isolated and damaged parts replaced. Another incident at Mauna Loa, causing a little more than 24 hours of lost data, was due to the suspension of a program that handles communications throughout the system. The cause is still unknown, but because other instruments at the observatory were disrupted simultaneously, the problem does not appear to be GONG specific. Sincere thanks to the on-site crew who contributed time to the GONG effort while tending to problems of their own.
The network instruments located at Learmonth and Mauna Loa suffered a combined down-time of about 47 hours due to a known intermittent problem. The data interrupt, which appears to be caused by a malfunctioning signal-generator card, is easily remedied by a reboot of the data acquisition system. Although the problem rarely occurs and is usually promptly corrected, a software patch may be implemented. The Exabyte tape drives continue to contribute their share to the number of lost images. About 14.5 hours of lost data accumulated at several sites due to the failure and replacement of problematic Exabyte drives.
The El Teide instrument was down for about a week during a scheduled preventative maintenance trip. This visit included all the routine maintenance tasks, as well as the replacement of several items that were compromising the backup capabilities of the instrument.
During the past quarter, month-long (36-day) velocity time series and power spectra were produced for GONG months 20 and 21 (ending 970625) with fill factors of 0.79 each. The fill factors for these two months were lower than those of the previous two months (0.92 and 0.89) but higher than the low of 0.73, which occurred during GONG month 13 (960712 through 960816).
The p-mode reprocessing campaign added GONG month 14 and the six month long (GONG months 14 through 19) time series, and power spectra were assembled.
During the current quarter, the project expects to reprocess months 12 and 13, to perform the original processing on months 22 and 23, and to produce a continuous ten-month long (one GONG year) time series and power spectra.
The project also began producing time series and power spectra from the intensity images. These products, generated from GONG intensity month 16, were produced using the new p-mode pipeline and archive. This was preceded by a test reduction of both the intensity and modulation images from month 16.
The Field Tape Reader (FTR) (the subsystem that receives the raw data cartridges from the observing sites) processed 94 cartridges (90 during the previous quarter) containing 554 (563) site-days from the seven instruments. 391 (378) site-days were calibrated.
During the past quarter, the DSDS serviced 14 (11 during the previous quarter) data distribution requests for 10,225 (3,286) files totaling 2.1 (0.9) Gigabytes of data. Each of these data requests was filled during the day on which the request was received. The DSDS performed 2,023 (1,532) data cartridge transactions (library check-ins and check-outs) in response to requests from the data reduction pipeline and other internal operations.
Substantial progress has been made in the testing and implementation of advanced spectral analysis methods for helioseismic data. Rudi Komm has been working with both multi-taper and wavelet denoising methods and has found that applying both techniques in tandem to the GONG data substantially improves the quality of the estimated oscillation parameters. Specifically, the number of good converged fits increases by 50%, and the estimated variances decrease by a factor of 10. There is also no evidence that any systematic bias is introduced into the results by these methods. This is all good news, and a full fit of GONG month 16 is now underway with the new processing.
I am delighted to welcome Pat Eliason, who took on the mantle of authority as GONG Project Manager on 1 September. She has worked in planetary and earth sciences in the past, as well as bossing a construction crew or two, so ought to be able to handle our rowdy bunch!
Ron Kroll, who has served as acting operations manager since mid-June, officially donned the manager's cap and has taken over network control.
Roy Tucker has also just joined GONG to push forward the proof of concept 1024² camera system. "Should be a piece of cake" is the sort of approach we are all pleased to have on board.
Winifred Williams left the project on 31 October. Winifred has made substantial contributions in the areas of the spherical harmonic transform and the merging algorithm. We wish her well.
We are also happy to announce the arrival of Rachel Howe, who joined the project on 1 November.
Guillermo Montijo, GONG's lead electronics technician throughout the deployment and initial operations phase, has moved over to work with the "dark side" here in NOAO. We are glad he will be close at hand and his ever cheery smile will still be part of the organization.